OpenStudy (anonymous):

Abdul, while driving to school, computes the average speed for his trip to be 20 km/hr. On his return trip along the same route, there is less traffic and the average speed is 40km/hr. What is the average sped for Abdul's trip?

OpenStudy (anonymous):

You need to find the total distance and the total time. The total distance is two times the distance to school, 2d.\[s=d/t=>t=d/s\]Where s is speed, d is distance and t is time. So the time taken to get to school is:\[d/s _{1}\]and the time taken to get back is:\[d/s _{2}\]so the total time is:\[d/s _{1}{+}d/s _{2}=d(1/s _{1}+1/s _{2})\]So, as speed = distance/time, divide the total distance by the total time:\[2d/d(1/s _{1}+1/s _{2})=2/(1/s _{1}+1/s _{2})=2/(1/20+1/40)=80/3\approx26.7km/hr\]

OpenStudy (anonymous):

@bendt.......I didn't understand the last part of the solution 2(1/20+1/40)=80/3=26.7 how? cn u plz explain

OpenStudy (anonymous):

Do you mean you don't understand why\[2/(1/20+1/40)=80/3\]or you don't understand how I got that expression?

OpenStudy (anonymous):

why? 2(1/20+1/40)=80/3

OpenStudy (anonymous):

\[2(1/20+1/40)=2(2/40+1/40)=2/(3/40)=80/3\]